OpenStudy (anonymous):

If Jessie is 24 years younger than her mother and if the sum of their ages is 84, how old is Jessie?

OpenStudy (zarkon):

(x+24)+x=84

OpenStudy (anonymous):

its either 54 , 30 , or 28

OpenStudy (zarkon):

(x+24)+x=84 2x+24=84 2x=60 x=30

OpenStudy (anonymous):

x+(x+24)=84 2x=84-24 x=30